OpenStudy (anonymous):

At the city museum, child admission is $5.30 and adult admission is $8.80. On Wednesday, 154 tickets were sold for a total sales of $1050.70. How many adult tickets were sold that day?

alones (alones):

you have any choices??

OpenStudy (anonymous):

what do you mean by that?

alones (alones):

i mean 75

OpenStudy (anonymous):

the answer is 67 tickets but i don't know how to solve that?

alones (alones):

i did like this so p+q=163 5.3p+9.1q=1198.3 53p+91q=11983 53p+53q=8639 38q=3344 q=88 p=75

alones (alones):

@SnuggieLad may you help this user...

OpenStudy (jiteshmeghwal9):

let the no adult tickets sld be x then child tickets sold=154-x 8.80x+5.30(154-x)=1050.70 solve this eqn for x

OpenStudy (anonymous):

thank you

OpenStudy (jiteshmeghwal9):

yw